7. Consider the following Demand-Supply Equations


Qd =30-2P

Qs =6+2P

i) What is equilibrium?

ii) Find the equilibrium price and quantity. Also draw the graph.


1
Expert's answer
2022-02-27T20:12:02-0500
  1. Economic equilibrium is a condition or state in which economic forces are balanced.
  2. Given that Qd=30-2P and Qs=6+2P

"E=Q_d=Q_s"

"30-2P=6+2P"

"4P=24"

"P^*=6"

"Q^*=6+2(6)=18"

Equilibrium price and quantity is 6 and 18 respectively.